OpenStudy (anonymous):

help me plzzz Solve for y. y/4-(y-1)/2=>3

OpenStudy (anonymous):

Okay, so the first thing you need to do is make the bottom denominator of (y/4)-(y/1) the same. This basically turns into (y/4)-(4y/4). From there you get y-4y/4 or -3y/4. With that now you need to divide -3y/4 by 2/1 (or 2). When you divide two numbers, you need to flip the second one and change the sign to multiplication, so now the equation becomes -3y/4*1/2.

OpenStudy (anonymous):

Now that you have -3y/4*1/2, you multiply and get \[-3y/8 \ge 3\]. If you multiply 8 on both sides you get \[-3y \ge 24\] When you divide negative numbers in an equation such as this, the greater than sign is flipped, therefore leading the answer to be \[y \le -8\] Hope this helped!